OpenStudy (anonymous):

17. A team of divers assembles at an elevation of -10ft relative to the surface of the water. Then the team dives at a rate of -50ft/min. Write a rule that represents the team's depth d as a function of time t. What is the team's depth after 3 mins?

OpenStudy (johnweldon1993):

Well depth here...starts off at -10ft...since they are already at a level below the surface of the water....so you'll have a -10 in your equation and the team dives at a rate of -50ft/min ...this can be represented by -50t so altogether you have D (depth) = -50t - 10 And to solve for the depth after 3 minutes....you just plug in (3) for 't' :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (johnweldon1993):

anytime :)

OpenStudy (anonymous):

Do you think you could help me with one more problem?

OpenStudy (johnweldon1993):

if I'm able to ...then yes :)

OpenStudy (anonymous):

Write a function rule for the area of a rectangle with a length 2ft less than 3 times its width. What is the area of the rectangle when its width is 2ft?

OpenStudy (johnweldon1993):

Okay, well the area of a rectangle...is Length Times Width right?

OpenStudy (anonymous):

right

OpenStudy (johnweldon1993):

And you know that the length...is 2ft less than 3 times the width...so L = 3w - 2 Now it GIVES you a width (2ft) so you have All you need to solve for area A = L * W A = (3(2) - 2) * 2 A = (6 - 2) * 2 A = 4 * 2 A = 8 Make sense? :)

OpenStudy (johnweldon1993):

AND, if you don't want to do all that equation stuff here....you could think about it Width = 2 Length is 2 less that 3 Times 2 so length is 2 less that 6 Length is 4 Area = Length * Width Area = 4 * 2 Area = 8 :)

OpenStudy (anonymous):

Okay I get it now Thank you very much :)

OpenStudy (johnweldon1993):

like I said before..anytime! :)